# Powering In-Database Dynamic Model Slicing for Structured Data Analytics

### v1

@@ 5. In-Database Model Inference
In this section, we present our in-database model inference system INDICES. We use PostgreSQL (Yu and Chen, 1995) as our underlying database system. By exploiting Postgres extension and User-Defined-Functions (UDFs), we seamlessly incorporate the SQL-aware dynamic model slicing technique LEADS onto PostgreSQL to enable in-database model inference.
The typical model inference pipeline consists of four stages: model loading, data retrieval, data preprocessing, and inference. A naive way to support LEADS is to decouple the database system and the inference system. That is, analysts retrieve data from the database via SQL query, preprocess the data, and perform inference in a dedicated inference system. However, this decoupled solution presents three drawbacks. First, moving the data out from the database can expose it to potential security risks and may not align with compliance standards. Second, it is troublesome for the users to maintain two separate systems with complicated data analytics workflow. For traditional data analysts accustomed to SQL queries, learning an additional inference framework represents an extra burden. Third, moving data from the database to the inference system may incur additional overhead and latency, particularly when dealing with large datasets being moved across the network to an external inference system. We conduct a profiling experiment to evaluate the time usage breakdown of the decoupled solution using LEADS with the PyTorch (Imambi et al., 2021) runtime. As shown in Figure 4, data retrieval time occupies approximately 33% to 43% of the total inference time. This overhead primarily stems from the database connection, serialization, network communication, and deserialization for data movement. Therefore, we focus on creating an in-database inference system by integrating the inference procedure onto PostgreSQL via UDF to avoid transferring data out from the database and reduce data retrieval overhead.
Figure 4. The breakdown response time of the inference stages for the decoupled approach.
@@ 5.1. Inference UDF Design
By utilizing a UDF that encompasses all stages of the inference process, users can initiate queries by specifying parameters such as ‘TableName’ and ‘WHERE’ conditions, as illustrated in Figure 5. The UDF then retrieves relevant subdatasets from the database by applying the ‘WHERE’ condition on ‘TableName’. Subsequently, the UDF dynamically loads the trained model and performs the designated inference task. Upon completion, the prediction results are returned to the users. However, relying solely on Python in UDFs for the entire model inference pipeline remains suboptimal due to its inefficient data retrieval process. Therefore, we introduce three optimizations to improve the UDF inference efficiency.
@@
Efficient Execution Allocation. We utilize a multi-language strategy in UDFs, combining low-level languages like C or Rust for efficient data retrieval, and high-level languages like Python for model loading, data preprocessing, and inference with its extensive ML libraries, such as Pytorch and Sklearn. Rust is employed in our system due to its Postgres extension development library PGRX (pgr, [n.d.]) which helps access advanced low-level data retrieval APIs in Postgres like the Server Programming Interface (SPI) for faster data retrieval. However, even with this approach, there are two main challenges for efficient model inference: (1) data copying overhead, arises from different execution environments and data representations between RDBMS and the inference runtime. It necessitates extensive copying and conversations. (2) state initializing overhead, comes from repeatedly loading and releasing the deep learning model when handling an inference request. To mitigate these overheads, we further design memory sharing and state caching techniques in INDICES to improve inference efficiency.
@@
Memory Sharing. Due to the isolation between Python and Rust execution environments, data transfer between them requires two read-write operations: first, data is fetched from an RDBMS and stored in Rust’s environment, then it is transferred from Rust’s memory to Python’s memory. To mitigate data transfer inefficiencies, we leverage shared memory to bypass redundant read-write operations. Initially, data is filtered and retrieved within Rust’s environment using SPI. The data is then directly written to shared memory. This shared memory, allocated at the commencement of UDFs invocation, is accessible in both environments. Thus, the Python environment can directly access and extract data, eliminating the need for an additional copying step.
Figure 5. INDICES inference UDF execution.
Figure 6. INDICES workflow: model training and in-database inference.
@@
State Caching. In handling numerous inference requests, the frequent loading and releasing of the model during each inference execution incur significant overhead. To address it, we persistently cache the general model trained via the LEADS  technique at the PostgreSQL session-level and maintain a state cache for the utilized sliced model. Specifically, when our inference UDF encounters an SQL query with a new filter condition, it first checks the cache for the existing sliced model related to this condition. If the corresponding dedicated model is not found, a new model is derived from the general model and stored in the cache. To ensure efficient memory utilization and achieve constant time complexity for management, we adopt the least recently used (LRU) caching policy to manage cached sliced models.
@@ 5.2. System Workflow
Next, we present INDICES’ workflow, which comprises training and in-database inference phases, as shown in Figure 6.
Training Phase. In the training phase, we construct a general model based on tables in RDBMS following LEADS for the prediction task. We collect SQL query logs from the real-world database to help construct the training workload. The frequent filter conditions in these logs reveal the attributes and features that data analysts consider most relevant and significant. Using these queries, we can extract the corresponding subdatasets as the training dataset. Both the SQL queries and the selected subdatasets are preprocessed, vectorized, and fed into the general model for iterative training (Step 1 in Figure 6). Once this well-trained general model is prepared, it is serialized and saved as a state dictionary (Step 2). When the associated UDF is invoked, the model is loaded into Postgres and dynamically sliced based on the SQL query, allowing for the handling of online inference requests.
In-Database Inference Phase. In the inference phase, we integrate the inference process into Postgres by implementing a UDF through extension installation. The UDF, named indices_inference, offers a SQL interface for issuing inference queries using the following statement:
SELECT indices_inference(<tableName>, <taskName>, <filter>);
which accepts three arguments, the tableName refers to the table in RDBMS from which the subdataset is selected, the taskName represents the prediction target (e.g., click-through-rate, readmission-rate). The final argument filter denotes the propositional formula following the “WHERE” clause.
@@
When the query parser receives the inference query (Step 3), it initiates the execution of the UDF within the PostgreSQL UDF runtime environment (Step 4), as illustrated in Figure 5. Upon activation, the UDF performs four main tasks for an online SQL query. First, using the initial two parameters, it identifies the prediction target and determines the required general model. It checks whether the model is already cached; if not, it locates and loads the trained general model (Step 5). Second, the UDF customizes the model via LEADS technique mentioned in Section 4 based on the filter in the query and subsequently caches it in the Least Recently Used (LRU) cache. Third, it retrieves the relevant data based on the filter conditions specified in the query via the server programming interface (SPI) and writes the selected data to a shared memory (Step 6). The model inference stage then reads the data from the shared memory and executes model inference. Finally, the UDF creates a view based on the original table, incorporating a new column that holds the predictive results.
6. Experiments
Table 2. Evaluation of performance improvements with LEADS.
DNN CIN AFN ARMNet
Datasets Metric w/o w/ Imprv. w/o w/ Imprv. w/o w/ Imprv. w/o w/ Imprv.
In this section, we evaluate the effectiveness of LEADS  and efficiency of our in-database inference system, INDICES, using four real-world datasets. Particularly, we devise the experiments to answer the following three key research questions (RQs):
  • RQ1: Does the LEADS  technique improve the SQL-Aware predictive modeling task compared with the original base models?
  • RQ2: How effective is each component of LEADS in these prediction tasks?
  • RQ3: Does the INDICES system improve the inference efficiency compared with the traditional decoupled approach?
We report our findings with regards to the above questions respectively in Sections 6.2, 6.3, and 6.4.
@@ 6. Experiments
6.1. Experimental Setup
6.1.1. Datasets
We conduct experiments on four real-world datasets from the domains of finance, sociology, and healthcare. The statistics of the datasets are summarized in Table 1.
(1) Payment (Yeh and Lien, 2009; Yeh, 2016) consists of the profile of credit card clients and their past bill payments. The task is to predict whether the payment on a credit card will be in default in the next month.
(2) Credit (Huang et al., 2020; Anna Montoya, 2018) is gathered by Home Credit Group, focusing on the unbanked population. The task is to predict the repayment abilities of this population for better loan experience.
(4) Diabetes (Strack et al., 2014; Clore and Strack, 2014) contains ten years of clinical care at 130 US hospitals. Each tuple pertains to hospital records of patients diagnosed with diabetes, including details like medications and laboratory results. The task is to predict the patient’s readmission.
6.1.2. Workloads
In SQL-aware predictive modeling, there is currently no benchmark that fulfills the criteria of having both SQL queries and supervised data. Traditional OLAP benchmarks, like TPC-DS (Nambiar and Poess, 2006) and YCSB (Cooper et al., 2010), primarily focus on assessing query performance with complex operations such as JOIN and GROUPBY. However, they lack prediction tasks and labeled data. On the other hand, conventional datasets for evaluating deep learning algorithms do not incorporate OLAP queries. To bridge this gap, we opt to create synthetic inference queries as workloads based on deep learning datasets to evaluate LEADS  and INDICES.
Our workload generation method is outlined in Algorithm 1, which employs a random strategy to generate a workload that comprises a set of synthetic SQL queries, with each query retrieving a subset of the dataset for prediction. The procedure begins by randomly selecting a data tuple x from the dataset $D$ (Step 3). Then, a value $m$ is sampled from the range $[1,\min(max\_{col},M)]$ to determine the number of predicates in the SQL query, where $M$ is the number of attributes in $D$ (Step 4). $max\_{col}$ is a parameter that indicates the maximum number of predicates in any SQL query. Subsequently, $m$ attributes are randomly chosen from x, and the values of these selected attributes are collected to form a propositional formula for the generated SQL query (Steps 5-6). This query is put into the workload (Step 7). We repeat this process $N$ times to create a complete workload. In the experiments, we set $N$ to 30 and $max\_col$ to 3, and generate workloads for each dataset.
Table 3. Top-4 SQL queries in terms of AUC improvement due to LEADS.
query# no. of tuples (test/train) propositional formula in query
1 20/134 change = “No” && admission_type = 3
2 20/153 outpatient = 20 && metformin.pio = “Up”
3 55/451 glipizide = “Down”
4 61/481 diag_1 = “50”
6.1.3. Baseline Methods
We select four kinds of base models designed for structured data and enhance these models via the LEADS technique. We evaluate LEADS’s effectiveness by comparing the performance of these base models with and without the integration of LEADS. We introduce each base model as follows.
(1) DNN (Gardner and Dorling, 1998): it is a perceptron with multiple linear and activation layers, representing the most fundamental neural network.
(2) CIN (Lian et al., 2018): it is a convolutional layer-based neural network, which models higher-order feature interactions through compressed interaction with input embeddings.
Moreover, to evaluate the efficiency of our in-database inference system INDICES, we compare it with INDICES-decoupled, a variant of INDICES that follows the traditional inference approach mentioned in Section 5. For INDICES-decoupled, since data is retrieved from PostgreSQL through network communication based on psycopg, there is no data copy process between different execution environments as in INDICES. To ensure a fair comparison, we warm up both systems by caching the general model in advance.
Figure 7. AUC improvement of SQL queries listed in Table 3.
(a) Payment
(b) Credit
(c) Census
(d) Diabetes
@@ 6.1.3. Baseline Methods
6.1.4. Evaluation Metric
Notice that a workload contains a series of prediction queries, with each query representing a specific prediction task. We use the AUC (Area Under the ROC Curve) metric to evaluate the effectiveness of LEADS on a query, denoted by $q$. A higher value indicates a better performance. Then, we use two metrics to assess the overall performance of LEADS on the entire workload. The first is the average AUC value across all queries in the workload, denoted as $Workload\text{-}AUC$, which is calculated by:
[…]
where $N$ is the number of SQL queries in the workload $W$. The second is the lowest AUC value among all prediction queries in the workload, termed $Worst\text{-}AUC$, which is calculated as follows:
[…]
In fields like finance or healthcare, where prediction errors can result in significant losses, focusing on the lower bound of the performance is crucial. The Worst-AUC metric provides insights into the worst-case scenario, ensuring that the technique’s performance is reliable and does not lead to the worst decisions.
For model-level efficiency, we utilize the floating-point operations per second (FLOPs) metric to measure the computations during the inference phase of a workload. As for our system INDICES, we measure the performance using the end-to-end response time, which calculates the CPU time elapsed from the moment a user invokes an inference query to the moment when the user receives the prediction results.
(a) Payment
(b) Credit
Figure 9. Effects of $\alpha$-entmax and number of experts on accuracy.
6.1.5. Hyper-parameter Settings
For fair comparisons, we fix the feature embedding size at 10 and set the size of the hidden layer to 32 for all the base models. Given the ability to select multiple experts in LEADS, we reduce the hidden layer size of each expert to 16 for better efficiency. The depth of each base model is set to 3. For ARMNet, we specify the number of heads and the hidden size of the self-attention module as 8 and 16, respectively. The initial $alpha$ in $\alpha$-entmax is set to 2.5. The number of experts of LEADS is searched in 2~256 and fixed at 16. Also, the balance regularization factor $\lambda_{1}$ and the sparsity regularization factor $\lambda_{2}$ are searched in 1e-3~5e-2, and fixed at 1e-3. We perform a sensitivity analysis on these hyper-parameters and report the best results.
6.1.6. Training Details
Since there are no specific SQL queries in the training dataset to update the parameters of the gating network, we simulate a SQL query for each input data following Steps 3-6 in Algorithm1. Additionally, we adopt the Adam (Kingma and Ba, 2014) optimizer with a learning rate searched in 1e-3~0.1 and a batch size of 1024 for all base models and datasets. All the experiments are conducted on a server equipped with Xeon(R) Silver 4114 CPU@2.2GHz (10 cores), 256G memory, and GeForce RTX 3090 Ti. We implement all the models with PyTorch 1.6.0 with CUDA 10.2.
(a) Payment
(b) Credit
Figure 10. Effects of $\alpha$-entmax and number of experts on efficiency.
6.2. SQL-aware Predictions
To evaluate the efficacy of LEADS, we investigate the performance improvement of four base models after integrating with LEADS for given workloads.
The experimental results are summarized in Table 2. The main observation is that the prediction performance w.r.t. both Workload-AUC and Worst-AUC consistently improve when utilizing LEADS for all base models and all workloads. Further, we note that the most significant improvement is in the Worst-AUC metric. For instance, when using DNN as the base model, LEADS achieves improvements of 55.76% and 16.00% on the Credit and Diabetes datasets, respectively. The reason for the base model’s low performance could be significant variability or nuances in the instances of the retrieved subset that are not well-represented in the training data. As a consequence, the trained base model fails to provide accurate predictions for these instances.
To further analyze the Worst-AUC improvement, we perform a breakdown analysis on the Diabetes dataset with a DNN base model. Table 3 describes the top-4 SQL queries in terms of AUC improvement due to LEADS, and Figure 7 presents the respective AUC improvement of these queries. We observe that the AUC values under these SQL queries are less than Workload-AUC (see Table 2), the average AUC value of all queries in the workload. In addition, the number of training/testing tuples in the selected subset for each SQL query is very small. For instance, there are only 134 training tuples for SQL query#1, while the whole training dataset contains 101,766 tuples (see Table 1). Without sufficient examples, the base model cannot generalize effectively in these subdatasets, resulting in a misleading prediction. In our LEADS, the SQL-aware network leverages the propositional formula from SQL queries as meta-information to assist the general model in learning associated patterns within these subdatasets. Therefore, a base model handled by LEADS yields better performance for queries with limited numbers of related training samples.
We note that the propositional formulas of these four queries in the Diabetes dataset all revolve around the drug’s status. For instance, in query #2, glipizide=“Down” denotes that the drug “glipizide” has been prescribed and its dosage reduced. In healthcare records, the drug’s status is sparse, seen in only a few training samples. Despite its rarity, it significantly affects readmission rates, making it essential for analysts to gain insights into retrieving these subsets. This makes LEADS  valuable for improving prediction models and preventing performance collapse, especially in healthcare, where insights from specific subsets matter. Our experimental results demonstrate that LEADS effectively handles this scenario. The improvement results from dynamically integrating multiple expert models. When faced with challenging input samples, LEADS strategically allocates more experts for the sliced model, enhancing the network’s complexity and improving prediction capabilities.
(a) Payment
(b) Credit

### v2

@@ 4. In-Database Model Inference
This section presents the implementation details of the in-database inference extension, which integrates the LEADS technique into NeurDB to deliver more accurate predictions. The typical inference pipeline consists of four stages: (1) model loading, where the model is loaded into memory; (2) data retrieval, where subdatasets specified by SQL queries are fetched; (3) data preprocessing, which transforms the raw data from the RDBMS into the tensor format for model inputs; and (4) inference execution, where the model performs forward computation to generate predictions. A straightforward approach to support LEADS is to decouple the database and inference systems, referred to as the Inference-Decouple Strategy (IDS). In this approach, analysts retrieve subdatasets from the database, preprocess them, and run inference in a dedicated external inference system. However, it has several drawbacks. First, exporting data from the database introduces security risks and may violate compliance regulations. Second, managing two separate systems complicates the analytics workflow, increases operational complexity, and imposes additional learning requirements. Third, transferring large volumes of data from the database to an external system incurs significant overhead and latency. Figure 4 illustrates the time usage breakdown of IDS when running inference tasks. Notably, data retrieval accounts for around 40% of the total inference time primarily due to the overhead from database connections, data I/O, and network communication. To address these issues, we utilize the In-database Inference Strategy (IIS), in which the database performs inference and avoids transferring data. This approach not only enhances security and simplifies workflow, but also accelerates the inference process.
Figure 5. Design and execution of inference UDF.
Figure 6. Workflow of in-database inference extension.
@@ 4.1. Inference UDF Design
For the implementation, we design UDFs to encapsulate all stages of the inference process and integrate them into the database as an extension. As illustrated in Figure 5., users begin by initiating predictive queries, specifying parameters ‘TableName’ and ‘WHERE’ conditions. The UDF then retrieves relevant subdatasets by applying the ‘WHERE’ condition on specified table, dynamically loads the trained model, and performs the inference. Upon completion, the results are returned to the users. However, directly embedding the decoupled inference process into UDFs is suboptimal due to Python’s performance limitation. Therefore, we introduce three key optimizations to improve UDF execution efficiency.
@@
Efficient Execution Allocation. To optimize UDF performance, we adopt a multi-language strategy, combining the strengths of both Rust and Python. Rust is used for efficient data retrieval, leveraging the PostgreSQL extension development library PGRX (pgr, 2023). This allows us to utilize PostgreSQL’s built-in APIs, the Server Programming Interface (SPI) for high-performance data access. Meanwhile, Python is employed for data preprocessing and inference, benefiting from its compatibility with machine learning libraries like PyTorch and Scikit-learn. This combination ensures that we balance efficiency in data retrieval with the flexibility required for advanced data processing and machine learning tasks. However, two main challenges remain for efficient inference: (1) data copying overhead, caused by the need for extensive data transfers between different execution environments. (2) state initialization overhead, resulting from the frequent loading of predictive models for each inference task. To address these issues, we introduce memory sharing and state caching in our UDF designed to improve efficiency.
@@
Memory Sharing. Data transfer between different execution environments requires two read-write operations: fetching data from RDBMS into Rust, then moving it to Python. To avoid this inefficiency, we introduce shared memory. Data is filtered in Rust using SPI and written directly to the shared memory, which is allocated before UDF invocation and accessible in both execution environment. The shared memory allows data to be used in Python without additional copying, eliminating unnecessary overhead.
@@
State Caching. As shown in Figure 4, model loading takes up about 10% of the total inference time per request. With many requests, the overhead of repeatedly loading the model grows. To reduce it, we implement session-level caching for the general model and maintain a state cache for the frequently used sliced models. Specifically, when processing a SQL query with a new filter condition, the UDF first checks the state cache for an existing sliced model. If none is found, a new sliced model is derived from the cached general model. We employ a Least Recently Used (LRU) cahcing strategy to manage the model cache, ensuring efficient memory use.
@@ 4.2. System Workflow
In this section, we describe the workflow of the in-database inference extension, which comprises both training and in-database inference phases as shown in Figure 6.
Training Phase. In the training phase, a general model is constructed according to the table schemas and the predictive task SQL query logs from real-world scenarios are collected as the representative workload for training. These queries are utilized to retrieve the corresponding subdatasets, both of which are then preprocessed and fed into the general model for iterative training (Step 1 in Figure 6). Upon completion of training, the general model is serialized and stored as a state dictionary (Step 2). In response to the incoming SQL query, an associated UDF is invoked, and then the model is loaded into the database memory and sliced to handle various online inference requests.
In-Database Inference Phase. The inference process is encapsulated into a UDF, which is integrated into database through extension installation. The UDF, named infer, provides a SQL interface for inference queries, as shown in the statement below:
SELECT infer(<tableName>, <task>, <propositional formula>);
Here, ¡tableName¿ indicates the source table for the subdataset. ¡task¿ specifies the prediction target (such as click-through rate or readmission rate) and directs to a particular deep learning model. ¡propositional formula¿ defines the filter conditions following the WHERE clause. Upon receiving the inference query (Step 3 in Figure 6), the UDF is executed as illustrated in Figure 5 (Step 4). The UDF performs four main tasks for an online inference request: (1) The prediction target and model are specified by ¡tableName¿ and ¡task¿. If the model is not cached, the UDF will locate and load the general model (Step 5). (2) The UDF slices this model according to the ¡propositional formula¿ under the guidance of the LEADS technique. (3) Predictive data is retrieved via the SPI and written into shared memory (Step 6). (4) The sliced model is applied to the data for inference. Finally, the UDF returns a view containing the original data alongside a new predictive column.
5. Experiments
In this section, we evaluate the effectiveness of LEADS  and efficiency of our in-database inference extension, using five real-world datasets. We first introduce the datasets and experimental setup, and then design experiments and report findings to address four key research questions (RQs):
  • RQ1: Does the LEADS  technique improve the SQL-aware predictive modeling task compared to base models?
  • RQ2: How effective is each component in LEADS technique?
  • RQ3: Does our in-database inference extension improve the efficiency compared to traditional approaches?
  • RQ4: How effective is our extension in complex scenarios?
5.1. Experimental Setup
5.1.1. Datasets
We conduct experiments on five real-world datasets from the domains of finance, sociology, and healthcare. The statistics of the datasets are summarized in Table 1.
(1) Payment (Yeh and Lien, 2009; Yeh, 2016) consists of credit card clients’ profiles and their past bill payments. The task is to predict if a credit card payment will default next month.
(2) Credit (Huang et al., 2020; Anna Montoya and Kotek, 2018) is gathered by Home Credit Group, focusing on the unbanked population. The task is to predict the repayment abilities of this population for better loan experience.
(4) Diabetes (Strack et al., 2014; Clore and Beata, 2014) comprises ten years of clinical care at 130 US hospitals. Each tuple pertains to hospital records of patients diagnosed with diabetes, including details like medications and laboratory results. The task is to predict the patient’s readmission.
(5) Avazu (Steve Wang, 2014) collects data from a mobile platform to predict click-through rates on ads. The dataset comprises millions of records and 22 attribute fields, covering mobile application and device information, with a total of 1,544,250 unique features.
5.1.2. Workloads
@@
Avazu 40,428,967 17.2% 22 1,544,250
In SQL-aware predictive modeling, traditional OLAP benchmarks like TPC-DS (Nambiar and Poess, 2006) and TPC-H focus on evaluating query performance using operations such as JOIN and GROUPBY, which lack prediction tasks. In contrast, existing prediction datasets used for evaluating predictive modeling performance do not incorporate SQL queries. To address this deficiency, we develop a method to sample synthetic inference queries for the evaluation of SQL-aware predictive modeling.
Our method, detailed in Algorithm 1, employs a random strategy to create a set of synthetic SQL queries. The generation procedure is as follows: (1) randomly select a data tuple x from the dataset $D$ (Step 3); (2) sample a value $m$ as the number of predicates in the current query from the range $[1,\min(max\_{col},M)]$, where $max\_{col}$ is the maximum number of predicates allowed in SQL queries, and $M$ is the number of attributes in $D$ (Step 4); (3) sample $m$ attributes from x and use their values to construct a propositional formula for the SQL query (Steps 5-6); (4) add the generated SQL query to the workload (Step 7). This procedure is repeated $N$ to create a comprehensive workload. In experiments, we set $N$ to 50 and $max\_col$ to 3, and generate workloads on each dataset accordingly.
5.1.3. Baseline Methods
We select four kinds of base predictive models and enhance these models via the LEADS technique. We evaluate LEADS’s effectiveness by comparing the performance of these base models with and without the integration of LEADS. The base models are as follows
(1) DNN (Gardner and Dorling, 1998): contains fully-connected linear and activation layers, representing the most fundamental neural network.
(2) CIN (Lian et al., 2018): it models higher-order feature combinations through compressed interaction with input embeddings.
Additionally, to evaluate the efficiency of the in-database inference extension, we compare it with the traditional analytic approach IDS described in Section 4, in which data is retrieved out of the database through network communication based on psycopg, while no data is copied between execution environments.
@@ 5.1.3. Baseline Methods
Table 2. Evaluation of performance improvements with LEADS.
DNN CIN AFN ARMNet
Datasets Metric w/o w/ Imprv. w/o w/ Imprv. w/o w/ Imprv. w/o w/ Imprv.
Workload-AUC 0.7355 0.7424 +0.94% 0.7324 0.7443 +1.62% 0.7364 0.7424 +0.81% 0.7387 0.7440 +0.71%
Avazu Worst-AUC 0.4231 0.5562 +31.46% 0.4531 0.5625 +24.14% 0.5031 0.5594 +11.19% 0.4615 0.5625 +21.88%
5.1.4. Evaluation Metric
A workload consists of a set of inference queries, each representing a prediction request. We evaluate the effectiveness of LEADS on a single inference query, denoted by $q$, using the AUC (Area Under the ROC Curve) metric, where higher values indicates better performance. To assess LEADS’s overall performance across the entire workload, we use two metrics. The first is the average AUC for all queries, denoted as Workload-AUC:
(10) $Workload\text{-}AUC(W)=\frac{1}{N}\sum_{i=0}^{N}AUC(q_{i})$
where $N$ is the number of SQL queries, $q_{i}$ is the $i$-th query in the workload $W$. The second is the lowest AUC value among all queries, termed Worst-AUC and calculated as:
(11) $Worst\text{-}AUC(W)=Min\{AUC(q_{1}),AUC(q_{2}),\dots,AUC(q_{N})\}$
In fields like finance and healthcare where mistakes can lead to significant losses, it’s important to focus on the worst-case performance. The Worst-AUC reveals whether the technique performs reliably and helps avoid poor decisions. To assess model-level efficiency, we utilize the floating-point operations per second (FLOPs) to measure the computations during the inference phase. For system-level evaluation, we measure the performance using response time, which tracks the CPU time elapsed from when a user initiates a query to when the prediction results are received.
5.1.5. Hyper-parameter Settings
For fair comparisons, we set the feature embedding size to 10 and the hidden layer size to 32 for all base models. Since LEADS allows multiple experts, we reduce the hidden layer size of each expert to 16 for efficiency. For ARMNet, we set the self-attention module to 4 heads and a hidden size of 16. The initial $\alpha$ in sparsemax is 2.5. The number of experts in LEADS is selected between 2 and 256 and fixed at 16. The balance regularization factor $\lambda_{1}$ and sparsity factor $\lambda_{2}$ are chosen between 1e-3 and 5e-2, both fixed at 1e-3. We conduct a sensitivity analysis on these hyperparameters and report the best results.
5.1.6. Training Details
Since the training dataset lacks specific SQL queries to update the gating network parameters, we simulate a SQL query for each input following Steps 3-6 in Algorithm 1 . We use the Adam optimizer (Kingma and Ba, 2014) with a learning rate range of 1e-3 to 0.1 and a batch size of 1024 for all base models and datasets. Experiments are conducted on a server with a Xeon Silver 4114 CPU @ 2.2GHz (10 cores), 256GB of memory, and a GeForce RTX 3090 Ti. All models are implemented using PyTorch 1.6.0 with CUDA 10.2.
Figure 7. Top-4 SQL queries in terms of AUC improvement.
5.2. SQL-aware Predictions
To answer the question RQ1, we investigate the performance of four base models with LEADS. and the results are summarized in Table 2. The main observation is that the prediction performance w.r.t. Workload-AUC and Worst-AUC consistently improve when utilizing LEADS on base models Notably, the most significant improvement is observed in the Worst-AUC metric. For instance, when using DNN as the base model, LEADS achieves improvements of 55.76% and 16.00% on the Credit and Diabetes datasets, respectively. The reason for the base model’s low performance could be significant variability or nuances in the instances of the retrieved subset that are not well-represented in the training data. Consequently, the trained base model fails to provide accurate predictions.
To further analyze the Worst-AUC improvement, we perform a breakdown analysis on the Diabetes dataset with the DNN base model. Figure 7 presents the top-4 SQL queries with the highest AUC improvements from LEADS, along with their query profiles. We notice that these queries have lower AUC values compared to the overall Workload-AUC (see Table 2) and involve small subsets of the training data. For instance, query#1 uses only 0.13% of the training data, about 130 out of 101,766 tuples. With such limited data, the base model struggles to generalize, resulting in poor predictions. LEADS addresses this by leveraging the propositional formulas in SQL queries to help the general model identify patterns in these small subdatasets, enhancing performance on queries with few training samples. In Diabetes dataset, many queries involve the drug status. For example, query#3 indicates a reduced dosage of glipizide, which significantly impacts readmission rates. This demonstrates LEADS’s value in improving healthcare predictions.
(a) Payment.
(b) Credit.
@@ 5.2. SQL-aware Predictions
Table 3. Effects of the number of predicates on AUC.
Dataset #predicates 1 2 3 4 5 6
Credit #tuple 24375 9902 9900 9896 9856 5208
w/o LEADS 0.7432 0.7495 0.7495 0.7494 0.7483 0.7494
w/ LEADS 0.7522 0.7589 0.7591 0.7581 0.7577 0.7604
(a) Payment-AUC.
(b) Credit-AUC.
(c) Payment-FLOPs.
(d) Credit-FLOPs.
Figure 9. Effects of $\alpha$-entmax and number of experts.
We also investigate the impact of predicate numbers on predictive results, as it represents query complexity. In our experiment, we incrementally added up to 6 predicates to evaluate the prediction. As shown in Table 3, LEADS outperforms baseline methods across different levels of query complexity. Additionally, there is an upward trend in predictive performance with the increase of the predicate number, indicating that complex primitive queries provide richer meta-information and enhance prediction accuracy.
(a) Payment.
(b) Credit.